# Possible Causes of False General Relativity Violations in Gravitational Wave Observations

### v1

@@ 5.2 Waveform Systematics
Employing non-precessing, eccentric waveforms, some papers have claimed the evidence for eccentricity in observed GW signals (Romero-Shaw et al., 2020; Gayathri et al., 2022; Romero-Shaw et al., 2021; O’Shea & Kumar, 2021; Gupte et al., 2024). Although this is contentious (see discussion in Section 3.1.1), it points to the fact that effects of eccentricity are already relevant in current observations, and therefore already pose a difficulty when performing tests of GR. This will continue to be a problem, and may be further exacerbated, as observatories become more sensitive.
@@
Tidal signatures may be present in several observed neutron star binary mergers (e.g., Abbott et al., 2017a, 2020b), although a confident detection of tidal signature is yet to occur. While misspecification of tidal effects is unlikely to appear as a GR violation in current detectors, a clean tidal signature may be present in A+ observatories for dynamical tidal effects (Pratten et al., 2022), and XG detectors for linear tides (e.g., Hinderer et al., 2010; Pereira et al., 2022).
@@
The kick-induced effects are too small to be detected with the current GW detectors but could potentially be observed in XG era (Gerosa & Moore, 2016; Mahapatra et al., 2023). The XG detectors are expected to observe $\sim 4-5$ events per year for which these effects will be constrained to better than $\sim 10\%$ (Mahapatra et al., 2023).
@@
Tests of GR and the no-hair theorem are already performed using the ringdown of loud GW signals (e.g., Abbott et al., 2016a) where the challenges that arise with specifying the ringdown start time and avoiding overfitting to nonlinearities are already present. These challenges will only intensify as the ringdown signals become louder in future observatories (e.g., Thrane et al., 2017).
@@
Several events in the existing GWTC have strong evidence of higher-order modes due, e.g., to extreme mass ratios such as GW190412 (Abbott et al., 2020d) and GW190814 (Abbott et al., 2020e). There are several events that have evidence of spin precession, such as GW190521 (Miller et al., 2024) and GW200129 (Hannam et al. (2022), although see Payne et al. (2022); Macas et al. (2024)). It is therefore important to account for spin precession and higher-order modes in current analyses, and the inclusion of higher modes will become even more important as the sensitivity of observatories continues to improve.
@@
Displacement memory is too small to be detected in individual events with the sensitivities of current detectors (Lasky et al., 2016; Hübner et al., 2020, 2021; Grant & Nichols, 2023). A memory signal is expected to influence parameter estimation results in loud events with SNR greater than 60, expected during the A# era (Xu et al., 2024), implying at this stage memory needs to be properly accounted for in waveforms models. Memory will have a significant influence in XG observatories; for example, Cosmic Explorer is predicted to have 3 to 4 events per year where memory is detectable for an individual event (Grant & Nichols, 2023), amplifying the need to properly account for memory effects.
@@
If we consider NR simulations to be the ground truth, then current waveform calibration errors refer to systematic biases introduced because the waveform approximants do not exactly match the NR simulations. But even NR waveforms carry uncertainties associated with, e.g., resolution effects and finite radius extraction. Such waveform calibration errors on the order of a few percent in amplitude, and a couple of degrees in phase, are subdominant to stochastic noise processes for binary neutron star observations at approximately 100 Mpc in A+ observatories (Read, 2023). Waveform uncertaintes are currently smaller than this, implying they are not a potential source of bias for tests of GR. This is not necessarily true in the A# and XG era when even NR waveforms will not be sufficiently accurate for unbiased parameter estimation recovery (Pürrer & Haster, 2020; Hu & Veitch, 2022). This latter point motivates the continual need for more accurate NR simulations and waveform extraction methods, as well as waveform approximations.
@@ 5.3 Astrophysical Aspects
In current and future detectors like advanced LIGO and A+, the estimated rate of strong lensing events for binary neutron stars is approximately $0.1\%$, while for binary black holes it is expected to be around $0.2\%$. These figures are consistent across various studies (Smith et al., 2023; Ng et al., 2018; Xu et al., 2022). Following this, advanced LIGO is anticipated to detect approximately 0.1 lensing events per year, whereas A+ is projected to observe 1 event annually. However, with XG detectors, ${\cal O}(100)$ events could be detected per year. It is important to note that these rates serve as a lower bound for millilensing and microlensing, since they could occur together with strong lensing in events. Therefore, lensing effects will not be a significant issue only until XG era.
@@
[…]
@@
The problem of fortifying hierarchical tests of GR against population assumptions and modelling systematics will be ever-present. Statistical assumptions on how to combine the information from individual events require care, as they reflect implicit assumptions on the beyond-GR theory that is being tested (Zimmerman et al., 2019; Isi et al., 2019b). Incorrect prior assumptions on the astrophysical population can cause biases if the deviation parameters are correlated with individual source parameters. These biases can be mitigated by jointly inferring the astrophysical population when performing hierarchical tests of GR, or in the high-SNR limit of XG detectors if the degeneracies between source parameters and deviation parameters are not perfect (Payne et al., 2023). Effects due to the finite size of the catalog (Pacilio et al., 2023) or selection effects against large deviations (Magee et al., 2024) can also lead to biases in population constraints if not properly accounted for. Finally, waveform systematics (both due to missing physics and sub-optimal calibration) can accumulate in a population analysis and lead to infer false GR violations even if the biases are under control at the single-event level (Moore et al., 2021; Saini et al., 2023). This effect will be even more prominent when restricting the test to high-SNR events that can be routinely observed with XG detectors (Hu & Veitch, 2023).
@@ Acknowledgments
A. Gupta, N.K. Johnson-McDaniel, and P. Narayan are supported by NSF Grants No. AST-2205920 and PHY-2308887. K.G. Arun acknowledges support from the Department of Science and Technology and Science and Engineering Research Board (SERB) of India via the Swarnajayanti Fellowship Grant DST/SJF/PSA-01/2017-18 and the Core Research Grant CRG/2021/004565. K.G. Arun and B.S. Sathyaprakash acknowledge the support of the Indo-US Science and Technology Forum through the Indo-US Centre for Gravitational-Physics and Astronomy, grant IUSSTF/JC-142/2019. E. Barausse acknowledges support from the European Union’s H2020 ERC Consolidator Grant “GRavity from Astrophysical to Microscopic Scales” (Grant No. GRAMS-815673), the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”, and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855. L. Bernard acknowledges financial support from the ANR PRoGRAM project, grant ANR-21-CE31-0003-001 and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement no. 101007855. E. Berti and L. Reali are supported by NSF Grants No. AST-2006538, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grants No. 20-LPS20-0011 and 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation grant No. PGR01167. V. Cardoso is a Villum Investigator and a DNRF Chair, supported by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. V. Cardoso acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. V. Cardoso has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101007855 and No. 101131233. S.Y. Cheung, T. Clarke, N. Guttman, P.D. Lasky, L. Passenger, and H. Tong are supported by Australian Research Council (ARC) Centre of Excellence for Gravitational-Wave Discovery CE170100004 and CE230100016, Discovery Projects DP220101610 and DP230103088, and LIEF LE210100002. S. Datta acknowledges support from UVA Arts and Sciences Rising Scholars Fellowship. A. Dhani, I. Gupta, R. Kashyap and B.S. Sathyaprakash were supported in part by NSF Grants No. PHY-2207638, AST-2307147, PHY-2308886, and PHYS-2309064. B. Sathyaprakash also thanks the Aspen Center for Physics (ACP) for hospitality during the summer of 2022. J.M. Ezquiaga is supported by the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847523 INTERACTIONS, and by VILLUM FONDEN (grant no. 53101 and 37766). E. Maggio acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) - project number: 386119226. E. Maggio is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101107586. A. Maselli acknowledges financial support from MUR PRIN Grants No. 2022-Z9X4XS and 2020KB33TP. S. Tiwari is supported by the Swiss National Science Foundation Ambizione grant no. PZ00P2-202204. K. Yagi is supported by NSF Grant PHY-2207349, PHY-2309066, PHYS-2339969, and the Owens Family Foundation. N. Yunes is supported by the Simons Foundation through Award No. 896696, the NSF Grant No. PHY-2207650, and the NASA Award No. 80NSSC22K0806.

### v2

@@ 5.2 Waveform Systematics
Employing non-precessing, eccentric waveforms, some papers have claimed the evidence for eccentricity in observed GW signals [128, 138, 139, 129, 140]. Although this is contentious (see discussion in Section 3.1.1), it points to the fact that effects of eccentricity are already relevant in current observations, and therefore already pose a difficulty when performing tests of GR. This will continue to be a problem, and may be further exacerbated, as observatories become more sensitive.
@@
Tidal signatures may be present in several observed neutron star binary mergers (e.g., [62, 502]), although a confident detection of tidal signature is yet to occur. While misspecification of tidal effects is unlikely to appear as a GR violation in current detectors, a clean tidal signature may be present in A+ observatories for dynamical tidal effects [503], and XG detectors for linear tides (e.g., [504, 505]).
@@
The kick-induced effects are too small to be detected with the current GW detectors but could potentially be observed in XG era [235, 234]. The XG detectors are expected to observe $\sim 4-5$ events per year for which these effects will be constrained to better than $\sim 10\%$ [234].
@@
Tests of GR and the no-hair theorem are already performed using the ringdown of loud GW signals (e.g., [1]) where the challenges that arise with specifying the ringdown start time and avoiding overfitting to nonlinearities are already present. These challenges will only intensify as the ringdown signals become louder in future observatories (e.g., [286]).
@@
Several events in the existing GWTC have strong evidence of higher-order modes due, e.g., to large mass ratios such as GW190412 [506] and GW190814 [507]. There are several events that have evidence of spin precession, such as GW190521 [508] and GW200129 ([70], although see [71, 509]). It is therefore important to account for spin precession and higher-order modes in current analyses, and the inclusion of higher modes will become even more important as the sensitivity of observatories continues to improve.
@@
Displacement memory is too small to be detected in individual events with the sensitivities of current detectors [510, 511, 512, 334]. A memory signal is expected to influence parameter estimation results in loud events with SNR greater than 90, expected during the A# era [331], implying at this stage memory needs to be properly accounted for in waveforms models. Memory will have a significant influence in XG observatories; for example, Cosmic Explorer is predicted to have 3 to 4 events per year where memory is detectable for an individual event [334], amplifying the need to properly account for memory effects.
@@
If we consider NR simulations to be the ground truth, then current waveform calibration errors refer to systematic biases introduced because the waveform approximants do not exactly match the NR simulations. But even NR waveforms carry uncertainties associated with, e.g., resolution effects and finite radius extraction. Such waveform calibration errors on the order of a few percent in amplitude, and a couple of degrees in phase, are subdominant to stochastic noise processes for binary neutron star observations at approximately 100 Mpc in A+ observatories [164]. Waveform uncertaintes are currently smaller than this, implying they are not a potential source of bias for tests of GR. This is not necessarily true in the A# and XG era when even NR waveforms will not be sufficiently accurate for unbiased parameter estimation recovery [501, 337]. This latter point motivates the continual need for more accurate NR simulations and waveform extraction methods, as well as waveform approximations.
@@ 5.3 Astrophysical Aspects
In current and future detectors like advanced LIGO and A+, the estimated rate of strong lensing events for binary neutron stars is approximately $0.1\%$, while for binary black holes it is expected to be around $0.2\%$. These figures are consistent across various studies [513, 514, 515]. Following this, advanced LIGO is anticipated to detect approximately 0.1 lensing events per year, whereas A+ is projected to observe 1 event annually. However, with XG detectors, ${\cal O}(100)$ events could be detected per year. It is important to note that these rates serve as a lower bound for millilensing and microlensing, since they could occur together with strong lensing in events. Therefore, lensing effects will not be a significant issue only until XG era.
@@
Astrophysical environments in which one may anticipate binary systems merging (and which may leave an imprint on the GW signal) include thick ($\bar{\rho}\sim 10^{-8}\text{ g}/\text{cm}^{3}$) and thin ($\bar{\rho}\sim 0.1\text{ g}/\text{cm}^{3}$) accretion disks around active galactic nuclei [32], cold dark matter spikes ($\bar{\rho}\sim 10^{-6}\text{ g}/\text{cm}^{3}$) [405], superradiant-boson clouds ($\bar{\rho}\sim 0.1\text{ g}/\text{cm}^{3}$[417] and the dynamical fragmentation of massive stars ($\bar{\rho}\sim 10^{7}\text{ g}/\text{cm}^{3}$[394]. Santoro et al. [443] found no support for environmental effects in GWTC-1, and found the environmental density would need to be $\sim 20\,{\rm g/cm^{3}}$ to be observable. This likely does not correspond to any of the astrophysical environments mentioned previously. For advanced LIGO design sensitivity, they find that dynamical friction effects are detectable at $\bar{\rho}$ $\gtrsim 10\,{\rm g/cm^{3}}$ for a GW170817-like event, while the effect of collisionless accretion is only visible for densities 10-100 times greater. As there are no proposed environments with such densities, it is unlikely for environmental effects to be visible in advanced LIGO data. They find that XG observatories will be sensitive to environmental densities of $\sim 10^{-3}\,{\rm g/cm^{3}}$, which includes both thin accretion disks and superradiant clouds. It is therefore likely that environmental signatures will only become relevant for GR tests in XG and beyond.
@@
The problem of fortifying hierarchical tests of GR against population assumptions and modelling systematics will be ever-present. Statistical assumptions on how to combine the information from individual events require care, as they reflect implicit assumptions on the beyond-GR theory that is being tested [496, 241]. Incorrect prior assumptions on the astrophysical population can cause biases if the deviation parameters are correlated with individual source parameters. These biases can be mitigated by jointly inferring the astrophysical population when performing hierarchical tests of GR, or in the high-SNR limit of XG detectors if the degeneracies between source parameters and deviation parameters are not perfect [493]. Effects due to the finite size of the catalog [500] or selection effects against large deviations [494] can also lead to biases in population constraints if not properly accounted for. Finally, waveform systematics (both due to missing physics and sub-optimal calibration) can accumulate in a population analysis and lead to infer false GR violations even if the biases are under control at the single-event level [491, 136]. This effect will be even more prominent when restricting the test to high-SNR events that can be routinely observed with XG detectors [90].
@@ Acknowledgements
A. Gupta, N.K. Johnson-McDaniel, and P. Narayan are supported by NSF Grants No. AST-2205920 and PHY-2308887. K.G. Arun acknowledges support from the Department of Science and Technology and Science and Engineering Research Board (SERB) of India via the Swarnajayanti Fellowship Grant DST/SJF/PSA-01/2017-18 and the Core Research Grant CRG/2021/004565. K.G. Arun and B.S. Sathyaprakash acknowledge the support of the Indo-US Science and Technology Forum through the Indo-US Centre for Gravitational-Physics and Astronomy, grant IUSSTF/JC-142/2019. E. Barausse acknowledges support from the European Union’s H2020 ERC Consolidator Grant “GRavity from Astrophysical to Microscopic Scales” (Grant No. GRAMS-815673), the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”, and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855. L. Bernard acknowledges financial support from the ANR PRoGRAM project, grant ANR-21-CE31-0003-001 and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement no. 101007855. E. Berti and L. Reali are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation grant No. PGR01167. V. Cardoso is a Villum Investigator and a DNRF Chair, supported by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. V. Cardoso acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. V. Cardoso has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101007855 and No. 101131233. S.Y. Cheung, T. Clarke, N. Guttman, P.D. Lasky, L. Passenger, and H. Tong are supported by Australian Research Council (ARC) Centre of Excellence for Gravitational-Wave Discovery CE170100004 and CE230100016, Discovery Projects DP220101610 and DP230103088, and LIEF LE210100002. S. Datta acknowledges support from UVA Arts and Sciences Rising Scholars Fellowship. A. Dhani, I. Gupta, R. Kashyap and B.S. Sathyaprakash were supported in part by NSF Grants No. PHY-2207638, AST-2307147, PHY-2308886, and PHYS-2309064. B. Sathyaprakash also thanks the Aspen Center for Physics (ACP) for hospitality during the summer of 2022. J.M. Ezquiaga is supported by the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No. 847523 INTERACTIONS, and by VILLUM FONDEN (grant no. 53101 and 37766). E. Maggio acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) - project number: 386119226. E. Maggio is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101107586. A. Maselli acknowledges financial support from MUR PRIN Grants No. 2022-Z9X4XS and 2020KB33TP. S. Tiwari is supported by the Swiss National Science Foundation Ambizione grant no. PZ00P2-202204. K. Yagi is supported by NSF Grant PHY-2207349, PHY-2309066, PHYS-2339969, and the Owens Family Foundation. N. Yunes is supported by the Simons Foundation through Award No. 896696, the NSF Grant No. PHY-2207650, and the NASA Award No. 80NSSC22K0806.